When the roller derby car reaches the bottom of the hill, how fast is it traveling (point C)? (Using g= 10m/s/s and creat am LOL to help solve the problem)
1
Expert's answer
2020-06-22T11:12:16-0400

Solution.


Initially, the roller car had potential energy, which turned into kinetic and work against friction forces.

"E_p=W_f+E_k;"

"mgh=\mu mgx+\dfrac{mv^2}{2}," x is the distance at which the friction force acted.

"\dfrac{v^2}{2}=gh-\mu gx;"

"v^2=2(gh-\mu gx);"

"v=\sqrt{2(gh-\mu gx)}=\sqrt{2g(h-\mu x)};"

Answer: "v=\sqrt{2g(h-\mu x)}."